The longest side of a rectangle is six less than three times the shortest side. if the perimeter of the rectangle is 85, what is the area of the rectangle to the nearest hundredth?
Found 3 solutions by Jeetbhatt10th, ikleyn, MathTherapy:
Answer by Jeetbhatt10th(11)

Let width be x units
Thus,length=3x-6
Perimeter =length +width
85=3x-6+x
85=4x-6
91=4x
x=22.75
Length=68.25-6=62.25
Area=length×width=62.25×22.75
=1416.1875=1416.19 sq.units (answer)

Answer by ikleyn(52781)

.
The longest side of a rectangle is six less than three times the shortest side. if the perimeter of the rectangle is 85,
what is the area of the rectangle to the nearest hundredth?
~~~~~~~~~~~~~~~~~~~~~~~~

Let width be x units

Thus, length = 3x-6.

Perimeter = 2*(length + width)

85 = 2*(3x-6+x)

85 = 8x-12

97 = 8x

x = 12.125

Length = 3*12.125-6 = 30.375

Area = length*width = 30.375*12.125

= 368.296875 sq.units (answer)


-------------------
Dear @Jeetbhatt10th !

In this forum, we just have one person who makes errors in every second solved problem.

So, such position is NOT vacant at this forum.


You started recently, but it is just third time I see your solution incorrect.

Please be more attentive in future.


I wish you to get every success in this forum.



Answer by MathTherapy(10552)


The longest side of a rectangle is six less than three times the shortest side. if the perimeter of the rectangle is 85, what is the area of the rectangle to the nearest hundredth?
Correct answer: